The seat R Us factory produces chairs with four legs and stools with three legs. The style of seats and legs are all the same this year. If all the leftover legs were put with the leftover seats to make chairs, one seat would be left over without any legs. However if all the leftover legs were put on the leftover seats to make stools, one leg would be left. Hoe many more legs than seats are in the leftover inventory?
Answer by macston(5194)

.
S=number of seats; L=number of legs
.
4(S-1)=L
3S+1=L
4(S-1)=3S+1
4S-4=3S+1
S=5 There are 5 seats.
.
L=3S+1
L=3(5)+1
L=15+1
L=16 There are 16 legs.
.
How many more legs than seats?
16 legs-5 seats=11 more legs
ANSWER There are 11 more legs than seats in inventory.